Know Thyself: Basic Addressing Patterns at the English Linguistics Department

Andrea Ágnes Reményi

0 Abstract

An earlier study (Reményi 1991, 1994) quantitatively proved that the choice among addressing patterns is influenced by basic sociological variables in an order characteristic of a group, and hypothesized that the order indicates the importance of the variables in a group’s value scale. The present pilot study quantifies addressing data as reported by the staff of this volume’s publisher, the English Linguistics Department of ELTE, to test the above hypothesis, and finds the group using considerably more informal addressing (the indicator of a solidarity-based setting), and being less influenced by acquired^{1} sociological variables (the indicators of a power-supporting situation) than the group studied earlier.

1 Introduction

What rules do we follow when addressing one another? Why am I on informal terms with X, and on formal terms with Y? And why do you do it just the other way round? Do we all follow the same norm at all?

The choices in the Hungarian system of address on all verbal, nominal and pronominal forms^{2} where reference is made to the listener(s) is based on a binary distinction between formality versus informality, or deference versus familiarity. This binary basis present in several languages (but following different rules) is usually referred to as tu/vos usage (‘T/V’ from now on).

Brown & Gilman (1960/1972) developed a system explaining similar addressing rules in a few European languages the following way: the norm follows a ‘power semantic’ whenever irreciprocal addressing is widespread copying the basic asymmetrical relationship, power itself; the norm follows a ‘solidarity semantic’ whenever irreciprocal addressing is absent, the norm being influenced by other factors, not power.

138 Andrea Ágnes Reményi

What is the function of a system of address? “In an adult group […] collocutors are obliged to mutually agree if they accept each other as members of the same group or as members of different groups, i.e. whether they are solidary or not” by way of the basic binary system of T/V (Reményi 1994 : 104). The importance of this agreement in a dyad’s communication is indicated by the frequency of reminders:

— the first (and often the only) message they communicate on meeting each other (greeting), and the last message on taking leave (saying goodbye)

— a constant reminder in every 2nd person verb form

— pronominal (and often nominal) address are all based on that choice.

What makes the addressing norm power- or solidarity-based? The norm of the smaller or larger speech community is influenced by its value system: in a solidarity-based system the choice to be solidary or non-solidary is based on several individual (sex, age, education, status, etc.) and relational attributes (differences in sex, age, education, status, etc.) acknowledged as important. In a power-based system only power-related attributes are considered important.

The present-day Hungarian adult norm can be called solidaritybased, as irreciprocity is practically absent. The few exceptions offer explanations contrary to Brown & Gilman’s (cf. Reményi 1994 : 102; for the single case in the present database: see below). On the other hand, the power semantic can be detected on subtler levels in the Hungarian system of address:

(i) In nominal and pronominal address-forms irreciprocity is widespread in dyads with a power gap. For example, the boss addresses the secretary by her first name (often adding a diminutive) e.g. Katika, while the secretary addresses him by his family name+title, or title+title, e.g. Kovács úr, Südi doktor, tanár úr. The choice between ön and maga is often reported to depend the same way on ‘up-talk’ vs. ‘down-talk’. The present paper discusses the basic binary distinction only, appearing most frequently in verbal inflections (2nd person for T, 3rd person for V). Nominal and pronominal address forms, greetings or the third verbal pattern (the tetszik-type) are excluded from the analysis.

(ii) In hierarchically rigid working communities the basic choice of T or V in the dyad is mostly influenced by acquired sociological variables that are important from the aspect of power (education and status).

Addressing Patterns at ELD 139

I.e. the more rigid the hierarchy is, the more dyads will decide to be on T- or V-terms on the basis of their place in the hierarchy. The present paper seeks to investigate this problem.

2 Data collection

All but one of the Hungarian staff (faculty and the secretary) of the English Linguistics Department of the ELTE and, due to the frequency of contact, two secretaries from other departments (n=18; from now on ‘department’) were interviewed at the end of October, 1994. Each interview included the single question whether the informant was on T- or V-terms with each of the other informants and with students^{3} in general. Apart from the answer, most informants contributed stories about the topic, of which notes were made. Basic personal data (SEX,^{4} AGE, degree (SCHooling), status (RANK)) and data on the frequency of routine contact (DSPAT) were also collected and quantified (e.g. RANK: 6 for egyetemi tanár ‘full professor’, 5 for docens ‘associate professor’, 4 for adjunktus ‘lecturer’, etc.).

The basic unit of a quantitative analysis of T/V-forms and all related usage must not be the individual (as is the case with other variable linguistic phenomena), because usage cannot be predicted on that basis alone. The choice between patterns is negotiated by the dyad.^{5} Brown & Ford (1972 :

128) rightly emphasize the importance of treating this type of data as “truly relational”.

Dyads were formed (n=153) whose values for absolute (i.e. individualbased) and relational (i.e. dyad-based) sociological variables were computed, and data for the linguistic variable (T/V) were added.

The following relational sociological variables received values for dyads by numerical comparison of individual informants’ data: DSEX (difference in sex), DAGE (in age), DSCH (in education/degree), DRANK (in status) and DSPAT (in spatial distance, a variable that tries to capture the frequency of routine contact). DSPAT was originally to be excluded, as 0 (=same room) and 1 (=same department) are difficult to separate, due to the reorganization of the department and reallocation of some rooms. But a few informants mentioned the influence of sharing their room with somebody in deciding to choose T, therefore it was used in the analysis.

It was hypothesized that the higher value we found in the relational sociological variables, the higher the value would be for the dependent linguistic variable (T=1, V=2), e.g. a smaller AGE or RANK difference would induce a higher chance of T-usage, while a bigger AGE or RANK difference would induce a higher chance of V-usage.

Absolute sociological variables were to be formed on the basis of the dyadic unit, as well: the sum of the two values in the dyad was calculated for SUMAGE (dyad-members’ age), SUMSCH (education), SUMRANK (status) to measure the influence of individual-based variables (as opposed to relative values, i.e. differences). It was hypothesized that a lower value for AGE, and a higher value for SCH and RANK would induce a lower value for the linguistic variable, i.e. a higher chance of T-usage.

All answers are double-checked, as both parties of each dyad were asked. The importance of this type of knowledge in relation to a speaker’s all acquaintances is indicated by the fact that only 3 informants were uncertain in one case each; but their partners in the given dyads resolved this uncertainty. There were only 2 dyads with conflicting information, in one of them the partners decided to change to T-terms between the interviews. One informant was not asked; data was filled by using the partners’ information in each dyad.

3 Results

Results for the department are presented and compared to those found at the previously studied pharmaceutical company (‘company’ from now on) (source: Reményi 1991, 1994).

Out of the 152 dyads using reciprocal T/V-forms, 114 (75%) are on T- and 38 (25%) are on V-terms. The ratio at the company was 466 (41.6%) to 654 (58.4%), respectively. This means that the overall proportion of T/Vusage already is drastically different between the two communities: T-usage is almost twice as widespread at the department.

The influence of the variables was judged by calculating correlations^{6} between the linguistic variable and each sociological variable and rankordering the coefficients (in descending order), where the value of significance for the coefficients was also considered.

The r correlation coefficient (with values between 1 and −1) indicates the interdependence of two variables. At r=1 there is a positive linear relationship between the variables (i.e. the higher the value of the first variable is, the higher that of the other will be). At r=0 there is no relationship between the variables. At r=−1 there is a negative linear relationship between the variables (i.e. the higher the value of the first variable is, the lower that of the other will be).

The p value of significance (between 0 and 1) indicates the likelihood of chance in accessing results. For example, p<.05 means less than a 5 per cent likelihood that we obtained the result due to chance.

Company  Department
(n=1120)  T/V  (n=152)  T/V
DSEX  .7055**  DSEX  .4869**
DSPAT  .3189**  DAGE  .3560**
DAGE  .2961**  SUMAGE  .3117**
DSCH  .2871**  DSPAT  .3064**
SUMAGE  .2249** (DSCH  .2485*)
DRANK  .1975** (DRANK  .2218*)
(SUMSCH  .0525)  (SUMRANK  .1368)
(SUMRANK  .0402)  (SUMSCH  .0047)

Table 1. Basic correlations of T/V and the sociological variables at the company and the department (* p < .01 ** p < .001)

Similarly to the company, DSEX, the difference of sex showed the strongest correlation with T/V-usage, though the relationship had only a lower, medium strength at the department. All variables connected to ascribed characteristics (sex and age) had a medium weight, while acquired characteristics (education and status) had only a negligible influence.

In other words, the overall group norm prescribes to group members to take mostly sex differences, age differences and their or their collocutor’s age in consideration when deciding to be on V-terms, otherwise T is to be used. (DSPAT, spatial distance also had a small, but significant influence over the choice.)

On the other hand, SUMRANK and SUMSCH had no influence in either group, i.e. no difference can be found between people in their T/Vusage as far as their status or educational level are concerned.

Company

(n=406)  T/V
DAGE  .5636**
SUMAGE  .5328**
SUMSCH  .2829**
SUMRANK  .2430**
DSPAT  .2093**
(DRANK  .1027)
(DSCH  .0519)

Table 2. Correlations of T/V and the sociological variables for female-to-female dyads at the company (no variation at the department)

Subgroups broken down by sex, however, behaved very differently. No variation was found in the female group: T-usage is absolute among the sample’s 28 female-to-female dyads (100%), i.e. nothing can force females to be on V-terms within their own sex—they are universally solidary with all other women in the group. Why is this so? The women at the company were sensitive mostly to age in same-sex dyads: both to age differences and absolute age. Why are women at the department not sensitive at least to those?

Company  Department
(n=32)  (n=8)
m  49.6  44.4
s  11.8  10.8

Table 3. Means and standard deviations of age for females at the company and the department

While some difference was found both in the mean and the standard deviation of age between the female groups at the company and the department (see Table 3), the difference is not significant (F =1.1974, p>.05), provided that both samples are normally distributed (a disputable and untestable condition at a sample-size where n=8). On the other hand, there are big age differences at the company, where the extremes in the female group are 20 vs. 70, 74 and 81, while at the department all 8 females are in their 30s to 50s. Both small sample size and small absolute range in age may also result in the invariable T-usage among women.

Company  Department
(n=131)  T/V  (n=44)  T/V
DSCH  .5506**  DAGE  .5904**
DSPAT  .5318**  DRANK  .4980**
SUMSCH  .5212**  DSCH  .4323*
DAGE  .3096** (SUMAGE  .2097)
DRANK  .3004** (SUMRANK  .1338)
(SUMAGE  .2642*) (DSPAT  .0690)
(SUMRANK  .1674)  (SUMSCH  .0462)

Table 4. Correlations of T/V and the sociological variables for male-to-male dyads at the company and the department

While male-to-male dyads at the company were sensitive to both relative and absolute education, the 2 male-to-male dyads at the department on V-terms (5%) do not help us to find any generalization.

Company  Department
(n=538)  T/V  (n=80)  T/V
DSPAT  .2227**  DAGE  .4506**
(DRANK  .1233*)  SUMAGE  .4456**
(SUMAGE  .1179*)  DSPAT  .3936**
(DSCH  .1093*) (DRANK  .2888*)
(DAGE  .0826)  (DSCH  .2858*)
(SUMSCH  .0373)  (SUMRANK  .1871)
(SUMRANK  .0164)  (SUMSCH  .0144)

Table 5. Correlations of T/V and the sociological variables for intersex dyads at the company and the department

[A bar chart appears here in the original, which is replaced by the data it contains.]

TOTAL F-to-F M-to-M F-to-M
department  75%  100%  95%  55%
company  41.6% 84%  58%  8%

Figure 1. The distribution of T-usage by groups and subgroups at the department and the company

Intersex dyads at the company were sensitive almost exclusively to the very difference in sex: not more than 8 per cent were on T-terms (only spatial distance had some other, weak influence on the choice). Intersex dyads at the department seem to be less sensitive to this deepest gap: 55 per cent are still on T-terms. Age, however, also influences the choice: the higher the age difference is, the more possible it is that the dyad will be on V-terms (DAGE). And the older the dyad members are, the higher the probability is that they will use V (SUMAGE). (Due to the scarcity of V among same-sex dyads, the order of variables influencing intersex dyads resemble that influencing the whole group, i.e. intersex dyads contribute the most in characterizing V-usage at the department.)

4 Discussion

The main aim of this pilot study was to test the hypothesis whether the subjective feeling of a less rigid hierarchy^{7} at the department can be detected in a difference in addressing patterns, mostly the ordering of basic sociological variables as factors influencing them. But, to the author’s surprise, some simpler differences having a possible explanatory power were also found: the quantity of T-usage was drastically different in the department with almost twice as many dyads being on T-terms as at the company. In compliance with the expectations, subtler, qualitative differences were also found, but at this point it is impossible to decide how and which of these differences connect to power relations.

The basic function of T-usage is to express solidarity. As T-usage is considerably higher at the department in general, and also in every subgroup broken down according to sexes (see Figure 1), the department seems to be a group more closely-knit by solidarity. (One source of more V-usage, i.e. emphasis on non-solidarity at the company is that two separate units, the ‘office’ and the ‘lab’, were studied there.)

Same-sex dyads behave very similarly at the department: T-usage is almost categorical in them. Those at the company seemed to follow different norms: female dyads were sensitive only to ascribed variables, while male dyads were also influenced by acquired ones. The main dividing line, the deepest trench, however, is between the sexes in both groups, as far as communication is concerned: T-usage is considerably lower in intersex dyads than in either of the same-sex subgroups.

On the other hand, the department differs basically from the company: ascribed variables influence both overall group and subgroup behaviour almost exclusively at the department — a phenomenon that has been proved to indicate a solidarity-supporting norm. (Due to scarce data, male-to-male dyads are excluded from this conclusion.) As opposed to this, the community at the company as a whole is more influenced by acquired variables, the indicators of a power-supporting setting. Thus, the statement that the norm at the department is solidarity based is supported both by the quantity and the quality of T/T-usage there.

What causes the department to define itself on the basis of a closelyknit solidarity? Why do the staff define most colleagues as in-group by expressing solidarity with T-usage? The answer probably is that while at the company no specific out-group forces the group to define its precise boundaries, at the university the huge mass of students forces the department to do so. The dividing line is defined even wider between faculty and students than between the sexes. While all members of the department employ T to define colleagues as in-group, not all of them employ V to define students as out-group. Informants told stories about having been forced by students, who wanted to define the lecturer as out-group by expressing non-solidarity, to use V by not returning T offered by the lecturer. The irreciprocal addressing was resolved by the lecturer giving in in each case mentioned, thus applying the adult-norm rule to avoid it. (A possible cause for irreciprocal usage in one of the dyads at the department can be a considerably high value for DAGE, DSCH and DRANK between them.)

Young students do not tolerate irreciprocity better because of their stubbornness. Some of them still have not been able to forget their child norm supporting power-related irreciprocity, and to fully acknowledge the adult norm suppressing its use.

notes

[1] The dichotomy ‘ascribed’ (for e.g. sex and age) and ‘acquired’ (for e.g. education and status) is borrowed from Preston.

[2] Greetings are also part of the system of address.

[3] Dyads including students were eventually excluded from the analysis.

[4] Capitals stand for the studied variables. Possible values (and given values): T/V: 1–2, SEX: 1–2, AGE: 0–100 (23–72), SCH: 1–6 (2–6), RANK: 1–6 (2–6), DSPAT: 0–2.

[5] For an explanation on negotiating address-usage as part of a language game, see Kiefer (1980).

[6] In this pilot study the small database did not make factor analysis (e.g.VARBRUL) viable.

[7] An (objective or subjective) sociological test developed to measure hierarchical rigidity is to be found and axiomatically used to prove that addressing patterns are a linguistic device to indicate it.

references

Brown, R. and M. Ford (1961/1972) ‘Address in American English’. Journal of Abnormal and Social Psychology 62. 375–385. Citation from Laver & Hutcheson (1972 : 128–145).

Brown, R. and A. Gilman (1960/1972) ‘The pronouns of power and solidarity’. In Sebeok, T. A., ed. (1960) Style in language. Cambridge, MA: MIT Press. 253–276. Citation from Laver & Hutcheson (1972 : 103–127).

Kiefer, F. (1980) ‘Greetings as language games’. Journal of Pragmatics 4. 147–155.

Laver, J. and S. Hutcheson, eds. (1972) Communication in Face to Face Interaction. Harmondsworth, Middlesex: Penguin.

Preston, D. R. (1989) Sociolinguistics and Second Language Acquisition. Oxford: Blackwell.

Reményi, A. Á. (1991) ‘Influences and values: A sociolinguistic study in the Hungarian system of address’. MA dissertation, Reading, England.

Reményi, A. Á. (1994) ‘Influences and values: A sociolinguistic study in the Hungarian system of address’. Studies in Applied Linguistics 1. 85–109.